Cousin and i tried working on this at 2pm an no luck.
write each equation of each hyperbola.
vertices (5,0) and (-5,0); focus (7,0)
We started by graphing the points, its Horizontal so we came to the conclusion that its (x-h)^2/ a^2 - (y-k)^2/ b^2 = 1

Answer by lwsshak3(11628)

write each equation of each hyperbola.
vertices (5,0) and (-5,0); focus (7,0)
**
Equation of a hyperbola with horizontal transverse axis is of the standard form:
(x-h)^2/a^2-(y-k)^2/b^2=1, with (h,k)=(x,y) coordinates of center.
for given equation:
center: (0,0)
length of transverse axis=10=2a
a=5
a^2=25
.
c=7
c^2=49
..
c^2=a^2+b^2
b^2=c^2-a^2=49-25=24
..
Equation of given hyperbola:
(x-h)^2/a^2-(y-k)^2/b^2=1
(x-0)^2/25-(y-0)^2/24=1
x^2/25-y^2/24=1